# The Oracle Complexity of Simplex-based Matrix Games: Linear Separability and Nash Equilibria

### v2

@@ 6 Discussion
As for simplex/simplex and $\ell_{1}$/simplex games, we prove an $\tilde{\Omega}(\epsilon^{-2/5})$ iteration lower bound for computing an $\epsilon$-suboptimal solution, for any algorithm using a two-sided oracle, implying the same lower bound for computing an $\epsilon$-approximate Nash equilibrium. This is an exponential improvement over the previously known lower bound for this task due to Hadiji et al. (2024), leaving open a $\tilde{O}(\epsilon^{-3/5})$ gap compared to the best known upper bound of $\tilde{O}(\epsilon^{-1})$.
@@ Acknowledgments
=$\stackrel{{\scriptstyle(1)}}{{=}}~{}\sum_{j=1}^{T}(v_{j-1,l}-v_{j,l})^{2}+v_{T,l}^{2}~{}\stackrel{{\scriptstyle(2)}}{{=}}~{}\sum_{j=1}^{T}2(v_{j-1,l}^{2}+v_{j,l}^{2})+v_{T,l}^{2}$
@@
Assuming $n\geq T$ (without loss of generality, since otherwise the probability lower bound in the lemma statement is less than 00), this can be loosely upper bounded by
@@
=$\stackrel{{\scriptstyle(5)}}{{=}}~{}\frac{\beta}{4}\left(\frac{4\alpha}{\beta}\mathbf{u}_{1}^{\top}\mathbf{w}-\sqrt{\sum_{j=1}^{T-1}(\mathbf{u}_{j}^{\top}\mathbf{w}-\mathbf{u}_{j+1}^{\top}\mathbf{w})^{2}+(\mathbf{u}_{T}^{\top}\mathbf{w})^{2}}\right)~{},$ (12)
@@
[…]
@@
$\max_{l\in[n]}\|A_{l}\|\leq 1~{},~{}~{}~{}\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}_{T+1}\leq 0~{},\text{~{}~{}~{}and~{}~{}~{}}\max_{\mathbf{w}:\|\mathbf{w}\|\leq 1}\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}~{}\geq~{}\frac{\alpha}{\sqrt{T+1}}~{}\geq\frac{1}{c^{\prime}T\sqrt{T\log(n)}}$
@@ Appendix B Missing proofs from Section 5
Consider $\mathbf{w}=\frac{1}{\sum_{t=1}^{T+1}\|\mathbf{u}_{t}\|_{1}}\sum_{t=1}^{T+1}\mathbf{u}_{t}$, which by construction satisfies $\|\mathbf{w}\|_{1}\leq 1$. Since $\mathbf{u}_{1},\ldots,\mathbf{u}_{T+1}$ are orthonormal, we have
$\left(\sum_{t=1}^{T+1}\|\mathbf{u}_{t}\|_{1}\right)A\mathbf{w}~{}$ $=~{}\left(\sum_{j=1}^{T}(\mathbf{v}_{j-1}-\mathbf{v}_{j})\mathbf{u}_{j}^{\top}+\mathbf{v}_{T}\mathbf{u}_{T+1}^{\top}\right)\left(\sum_{t=1}^{T+1}\mathbf{u}_{t}\right)$
@@
Moreover, recalling the construction of $\mathbf{u}_{t}$ in Construction 2, we have by the second half of Lemma C.1 that
$\Pr\left(\|\mathbf{u}_{t}\|_{\infty}\geq\frac{z}{\sqrt{d}}~{}\middle|~{}\boldsymbol{\xi}^{\prime}_{1},\ldots,\boldsymbol{\xi}^{\prime}_{t-1}\right)~{}\leq~{}2d\exp\left(-\frac{z^{2}}{8}\right)+\exp\left(-\frac{d}{48}\right)~{}.$
This holds for any realization of $\boldsymbol{\xi}^{\prime}_{1},\ldots,\boldsymbol{\xi}^{\prime}_{t-1}$, hence by a union bound
$\Pr\left(\max_{t\in[T+1],i}|u_{t,i}|\geq\frac{z}{\sqrt{d}}\right)~{}\leq~{}(T+1)\left(2d\exp\left(-\frac{z^{2}}{8}\right)+\exp\left(-\frac{d}{48}\right)\right)~{}.$
Equivalently, by equating $2d\exp(-z^{2}/8)$ to $\delta$, we have that for any $\delta\in(0,1)$, with probability at least $1-(T+1)\left(\delta+\exp(-d/48)\right)$,
$\max_{t\in[T+1]}\|\mathbf{u}_{t}\|_{\infty}\leq\sqrt{\frac{8\log(2d/\delta)}{d}}~{},$ (15)
in which case
$\sum_{t=1}^{T+1}\|\mathbf{u}_{t}\|_{1}~{}\leq~{}(T+1)d\sqrt{\frac{8\log(2d/\delta)}{d}}~{}=~{}(T+1)\sqrt{8d\log(2d/\delta)}~{}.$
Coarsely upper bounding $(T+1)$ by $2T$ and plugging into Eq. (14), it follows that with probability at least $1-(T+1)\left(\delta+\exp(-d/48)\right)\geq 1-2T(\delta+\exp(-d/48))$, each entry of $A\mathbf{w}$ is at least
$\frac{\alpha}{2T\sqrt{8d\log(2d/\delta)}}~{}.$
This holds for the $\mathbf{w}$ we have chosen, so the maximum of $\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}$ over all vectors $\mathbf{w}$ in the unit $\ell_{1}$ ball can only be larger, from which the lemma follows.
@@
Recalling the construction of $\mathbf{v}_{t}$ in Construction 1, and applying Lemma C.1 combined with a union bound, it follows that for any $z>0$,
$\Pr\left(\max_{t\in[T]}\|\mathbf{v}_{t}\|_{\infty}>z\right)~{}\leq~{}2Tn\exp\left(-\frac{z^{2}}{2\beta^{2}}\right)~{}.$
Equivalently, by choosing $z$ appropriately, we have that for any $\delta\in(0,1)$, with probability at least $1-\delta$,
$\max_{t\in[T]}\|\mathbf{v}_{t}\|_{\infty}~{}\leq~{}\beta\sqrt{2\log(2Tn/\delta)}~{}.$
Also, applying the same lemma with respect to $\mathbf{u}_{t}$ (as done in Eq. (15) above), we have that with probability at least $1-(T+1)\left(\delta+\exp(-d/48)\right)$,
$\max_{t\in[T+1]}\|\mathbf{u}_{t}\|_{\infty}\leq\sqrt{\frac{8\log(2d/\delta)}{d}}~{}.$
Assuming these two events hold (with probability at least $1-\delta-(T+1)\left(\delta+\exp(-d/48)\right)$, by a union bound), and recalling that $v_{0,l}=\alpha$ for all $l$ by construction, it holds for any $l\in[n],i\in[d]$ that
$|A_{l,i}|~{}$ $=~{}\left|\sum_{j=1}^{T}(v_{j-1,l}-v_{j,l})u_{j,i}+v_{T,l}u_{T+1,i}\right|$
$\leq~{}\sum_{j=1}^{T}|v_{j-1,l}-v_{j,l}|\cdot|u_{j,i}|+|v_{T,l}|\cdot|u_{T+1,i}|$
$\leq~{}\sqrt{\frac{8\log(2d/\delta)}{d}}\left(\sum_{j=1}^{T}|v_{j-1,l}-v_{j,l}|+|v_{T,l}|\right)$
$\leq~{}\sqrt{\frac{8\log(2d/\delta)}{d}}\left(\sum_{j=1}^{T}\left(|v_{j-1,l}+|v_{j,l}|\right)+|v_{T,l}|\right)$
$\leq~{}\sqrt{\frac{8\log(2d/\delta)}{d}}\left(\alpha+2T\beta\sqrt{2\log(2Tn/\delta)}\right)~{}.$
By the assumption stated in the lemma, this expression is at most $1$, hence the entries of $A$ are bounded in $[-1,+1]$ as required. All this holds with probability at least $1-\delta-(T+1)\left(\delta+\exp(-d/48)\right)$, which can be coarsely lower bounded by $1-3T\left(\delta+\exp(-d/48)\right)$.
@@
We construct the matrix $A$ as in Construction 2. Examining the conditions in Proposition 4.6, and Lemma 5.3 and Lemma 5.4, and ignoring log factors momentarily, we see that in order for all of them to be applicable, we must choose $\alpha\lesssim\frac{\beta}{\sqrt{T}}$ (Proposition 4.6), satisfy $\beta\lesssim\frac{\sqrt{d}}{T}$ and $\alpha\lesssim\sqrt{d}$ (Lemma 5.4), yet the game value guarantee is $\approx\frac{\alpha}{T\sqrt{d}}$ (Lemma 5.3). Thus, to make the game value $\frac{\alpha}{T\sqrt{d}}$ as large as possible, yet satisfying all of the constraints, we should choose $\beta\approx\frac{\sqrt{d}}{T}$ and $\alpha\approx\frac{\beta}{\sqrt{T}}\approx\sqrt{\frac{d}{T^{3}}}$, and get a margin guarantee of approximately $\frac{\alpha}{T\sqrt{d}}\approx\frac{1}{T^{2}\sqrt{T}}$.
A bit more formally, pick (say) $\delta=1/(50T)$ in both Proposition 4.6, Lemma 5.3 and Lemma 5.4, and choose
$\beta=\frac{\sqrt{d}}{cT\sqrt{\log(nT)\log(dT)}}$
for some sufficiently large universal constant $c$, as well as
$\alpha=\frac{\beta}{4\sqrt{T}}=\frac{\sqrt{d}}{4cT\sqrt{T\log(nT)\log(dT)}}~{}.$
@@
$\max_{\mathbf{w}:\|\mathbf{w}\|\leq 1}\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}~{}\geq~{}\frac{\alpha}{c^{\prime}T\sqrt{d\log(dT)}}~{}=~{}\frac{1}{4cc^{\prime}\log(dT)\sqrt{\log(nT)}}\cdot\frac{1}{T^{2}\sqrt{T}}~{},$
all with some positive probability. Hence, by the probabilistic method, a suitable fixed matrix $A$ satisfying all of the above must exist. Coarsely upper bounding $\log(dT)$ by $\log(d^{2})=2\log(d)$, and $\log(nT)$ by $\log(n^{2})=2\log(n)$, and plugging in the displayed equation above, the theorem follows.
Appendix C Auxiliary lemma

### v3

@@ 6 Discussion
As for simplex/simplex and $\ell_{1}$/simplex games, we prove an $\tilde{\Omega}(\epsilon^{-2/3})$ iteration lower bound for computing an $\epsilon$-suboptimal solution, for any algorithm using a two-sided oracle, implying the same lower bound for computing an $\epsilon$-approximate Nash equilibrium. This is an exponential improvement over the previously known lower bound for this task due to Hadiji et al. (2024), leaving open a $\tilde{O}(\epsilon^{-1/3})$ gap compared to the best known upper bound of $\tilde{O}(\epsilon^{-1})$.
@@ Acknowledgments
$\stackrel{{\scriptstyle(1)}}{{=}}~\sum_{j=1}^{T}(v_{j-1,l}-v_{j,l})^{2}+v_{T,l}^{2}~\stackrel{{\scriptstyle(2)}}{{=}}~\sum_{j=1}^{T}2(v_{j-1,l}^{2}+v_{j,l}^{2})+v_{T,l}^{2}$
@@
Assuming $n\geq T$ (without loss of generality, since otherwise the probability lower bound in the lemma statement is less than 0), this can be loosely upper bounded by
@@
$\stackrel{{\scriptstyle(5)}}{{=}}~\frac{\beta}{4}\left(\frac{4\alpha}{\beta}\mathbf{u}_{1}^{\top}\mathbf{w}-\sqrt{\sum_{j=1}^{T-1}(\mathbf{u}_{j}^{\top}\mathbf{w}-\mathbf{u}_{j+1}^{\top}\mathbf{w})^{2}+(\mathbf{u}_{T}^{\top}\mathbf{w})^{2}}\right)~,$ (12)
@@
$0$~ $<~\left(\delta\mathbf{e}_{1}^{\top}\mathbf{x}-\sqrt{\mathbf{x}^{\top}M\mathbf{x}}\right)\left(\delta\mathbf{e}_{1}^{\top}\mathbf{x}+\sqrt{\mathbf{x}^{\top}M\mathbf{x}}\right)$
@@
$\max_{l\in[n]}\|A_{l}\|\leq 1~,~~~\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}_{T+1}\leq 0~,\text{~~~and~~~}\max_{\mathbf{w}:\|\mathbf{w}\|\leq 1}\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}~\geq~\frac{\alpha}{\sqrt{T+1}}~\geq\frac{1}{c^{\prime}T^{3/2}\sqrt{\log(n)}}$
@@ Appendix B Missing proofs from Section 5
Consider $\mathbf{w}=\frac{1}{\|\sum_{t=1}^{T+1}\mathbf{u}_{t}\|_{1}}\sum_{t=1}^{T+1}\mathbf{u}_{t}$, which by construction satisfies $\|\mathbf{w}\|_{1}\leq 1$. Since $\mathbf{u}_{1},\ldots,\mathbf{u}_{T+1}$ are orthonormal, we have
$\left\|\sum_{t=1}^{T+1}\mathbf{u}_{t}\right\|_{1}\cdot A\mathbf{w}$~ $=~\left(\sum_{j=1}^{T}(\mathbf{v}_{j-1}-\mathbf{v}_{j})\mathbf{u}_{j}^{\top}+\mathbf{v}_{T}\mathbf{u}_{T+1}^{\top}\right)\left(\sum_{t=1}^{T+1}\mathbf{u}_{t}\right)$
@@
Moreover, it holds that
$\left\|\sum_{t=1}^{T+1}\mathbf{u}_{t}\right\|_{1}\leq\sqrt{d}\left\|\sum_{t=1}^{T+1}\mathbf{u}_{t}\right\|_{2}=\sqrt{d(T+1)}~,$
and so by Eq. (14), each entry of $A\mathbf{w}$ is at least $\alpha/\sqrt{d(T+1)}$. This holds for the $\mathbf{w}$ we have chosen, so the maximum of $\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}$ over all vectors $\mathbf{w}$ in the unit $\ell_{1}$ ball can only be larger, from which the lemma follows.
@@
For any $l\in[n],i\in[d]$ it holds that
$|A_{l,i}|=\left|\sum_{j=1}^{T}(v_{j-1,l}-v_{j,l})u_{j,i}+v_{T,l}u_{T+1,i}\right|\leq\left|\sum_{j=1}^{T}(v_{j-1,l}-v_{j,l})u_{j,i}\right|+|v_{T,l}|\cdot|u_{T+1,i}|~.$ (15)
We turn to bound the terms above. For $l\in[n],i\in[d]$, we denote $D^{j}_{l}:=v_{j-1,l}-v_{j,l}$ and $S_{l,i}:=\sum_{j=1}^{T}D^{j}_{l}u_{j,i}$. Note that conditioned on $\mathbf{v}_{0},\dots,\mathbf{v}_{T}$, each $u_{j,i}$ is a sub-Gaussian, mean-zero random variable with sub-Gaussian norm bounded by $O(1/d)$, and therefore $S_{l,i}$ has sub-Gaussian norm at most $O(\frac{1}{d})\cdot\sum_{j=1}^{T}(D_{l}^{j})^{2}$ (cf. Vershynin, 2018 for background on sub-Gaussian random variables). In other words, there exists a universal constant $C_{1}>0$ such that for any $z>0:$
$\Pr\left(|S_{l,i}|\geq z\mid\mathbf{v}_{0},\dots,\mathbf{v}_{T}\right)\leq 2\exp\left(-\frac{C_{1}dz^{2}}{\sum_{j=1}^{T}(D_{l}^{j})^{2}}\right)~.$
The numerical inequality $(a-b)^{2}\leq 2(a^{2}+b^{2})$ implies that $\sum_{j=1}^{T}(D^{j}_{l})^{2}\leq 4\sum_{j=0}^{T}v_{j,l}^{2}=4\alpha^{2}+4\sum_{j=1}^{T}v_{j,l}^{2}$, and further note that while $v_{1,l},\dots,v_{T,l}$ are not independent, it does hold that conditionally $v_{t,l}\mid v_{0,l},\dots v_{t-1,l}\sim\mathcal{N}(0,\sigma_{t,l}^{2})$ for some $\sigma_{t,l}\leq\beta$ since projecting a Gaussian cannot increase its variance. Therefore, Lemma C.2 ensures that with probability at least $1-\delta:~\sum_{j=1}^{T}v_{j,l}^{2}\leq 8\beta^{2}(T+\log(1/\delta))$, so via a union bound, we get that with probability at least $1-2\delta$ it holds for all $z>0$ that
$\Pr\left(|S_{l,i}|\geq z\right)\leq 2\exp\left(-\frac{C_{1}dz^{2}}{4\alpha^{2}+32\beta^{2}(T+\log(1/\delta))}\right)~.$
Equivalently, choosing $z$ by equating the above to $\delta$, slightly simplifying the expression and union bounding over $l\in[n]$ and $i\in[d]$, we see that for some sufficiently large absolute constant $C>0$ it holds with probability at least $1-3nd\delta$ that
$\left|\sum_{j=1}^{T}(v_{j-1,l}-v_{j,l})u_{j,i}\right|=|S_{l,i}|\leq C\sqrt{\frac{\log(1/\delta)}{d}}\left[\alpha+\beta(\sqrt{T}+\sqrt{\log(1/\delta)})\right]~,$ (16)
for all $l\in[n],i\in[d]$ simultaneously.
To bound the second term in Eq. (15), by applying Lemma C.1 (with appropriately chosen $z$) it follows that for any $\delta\in(0,1)$, with probability at least $1-\delta:$
$\|\mathbf{v}_{T}\|_{\infty}~\leq~\beta\sqrt{2\log(2n/\delta)}~.$ (17)
Also, applying the same lemma with respect to $\mathbf{u}_{T+1}$, we have that with probability at least $1-\left(\delta+\exp(-d/48)\right):$
$\|\mathbf{u}_{T+1}\|_{\infty}\leq\sqrt{\frac{8\log(2d/\delta)}{d}}~.$ (18)
Overall, under the desired events of Eqs. (16), (17) and (18), we see by Eq. (15) for all $l\in[n],i\in[d]:$
$|A_{l,i}|$ $\leq C\sqrt{\frac{\log(1/\delta)}{d}}[\alpha+\beta(\sqrt{T}+\sqrt{\log(1/\delta)})]+4\beta\sqrt{\frac{\log(2n/\delta)\log(2d/\delta)}{d}}$
$\leq C\frac{\log((n+d)/\delta)}{\sqrt{d}}\left(\alpha+\beta\sqrt{T}+1\right)~,$
the latter holding as long as $C$ is a sufficiently large absolute constant. For any $\delta\in(0,1)$ such that the expression above is at most $1$, which is equivalent to the assumption stated in the lemma, we see that $A\in[-1,1]^{n\times d}$ as required. All this holds with probability at least $1-3nd\delta-2\delta-\exp(-d/48)$ by a union bound, which can be coarsely lower bounded by $1-5nd\delta-\exp(-d/48)$.
@@
We construct the matrix $A$ as in Construction 2. Examining the conditions in Proposition 4.6, and Lemma 5.3 and Lemma 5.4, and ignoring log factors momentarily, we see that in order for all of them to be applicable, we must choose $\alpha\lesssim\frac{\beta}{\sqrt{T}}$ (Proposition 4.6), satisfy $\beta\lesssim\sqrt{d/T}$ and $\alpha\lesssim\sqrt{d}$ (Lemma 5.4), yet the game value guarantee is $\approx\frac{\alpha}{\sqrt{Td}}$ (Lemma 5.3). Thus, to make the game value $\frac{\alpha}{\sqrt{Td}}$ as large as possible, yet satisfying all of the constraints, we should choose $\beta\approx\sqrt{d/T}$ and $\alpha\approx\beta/\sqrt{T}\approx\sqrt{d}/T$, and get a margin guarantee of approximately $\frac{\alpha}{\sqrt{Td}}\approx\frac{1}{T^{3/2}}$.
A bit more formally, pick (say) $\delta=1/(50nd)$ in both Proposition 4.6, Lemma 5.3 and Lemma 5.4, and choose
$\beta=\frac{\sqrt{d}}{3C\sqrt{T}\log(50nd(n+d))}$
for the sufficiently large universal constant $C$ given by Lemma 5.4, as well as
$\alpha=\frac{\beta}{4\sqrt{T}}=\frac{\sqrt{d}}{12CT\log(50nd(n+d))}~.$
@@
$\max_{\mathbf{w}:\|\mathbf{w}\|_{1}\leq 1}\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}\geq\frac{\alpha}{\sqrt{d(T+1)}}=\frac{\beta}{4\sqrt{T}}=\frac{1}{12C\log(50nd(n+d))}\cdot\frac{1}{T^{3/2}}$
all with some positive probability. Hence, by the probabilistic method, a suitable fixed matrix $A$ satisfying all of the above must exist. Coarsely upper bounding $\log(50nd(n+d))\leq\log(100n^{2}d^{2})=2\log(10nd)$, and plugging in the displayed equation above, the theorem follows.
Appendix C Auxiliary lemmas
@@
Lemma C.2.
Let $\xi_{1},\dots,\xi_{T}$ be a sequence of random variables such that for all $t\in[T]:~\xi_{t}\mid\xi_{1},\dots,\xi_{t-1}\sim\mathcal{N}(0,\sigma_{t}^{2})$, and further assume that $\sigma_{t}^{2}\leq\beta^{2}$ for all $t$. Then with probability at least $1-\delta:~\sum_{t=1}^{T}\xi_{t}^{2}\leq 8\beta^{2}(T+\log(1/\delta))$.
Proof.
Denote $S:=\sum_{t=1}^{T}\xi_{t}^{2},~V:=\mathbb{E}[S]=\sum_{t=1}^{T}\sigma_{t}^{2}\leq\beta^{2}T$ (by the law of total expectation). Fix $\lambda\in[0,1/4\beta^{2}]$ and let
$M_{t}:=\exp(\lambda\sum_{i=1}^{t}\xi_{i}^{2})\prod_{i=1}^{t}\sqrt{1-2\lambda\sigma_{i}^{2}}~,$
where we also let $M_{0}:=1$. Then
$\mathbb{E}[M_{t}\mid\xi_{1},\dots,\xi_{t-1}]=\exp(\lambda\sum_{i=1}^{t}\xi_{i}^{2})\prod_{i=1}^{t}\sqrt{1-2\lambda\sigma_{i}^{2}}\cdot\mathbb{E}[e^{\lambda\xi_{t}^{2}}\mid\xi_{1},\dots,\xi_{t-1}]\sqrt{1-2\lambda\sigma_{t}^{2}}=M_{t-1}~,$
thus $(M_{t})_{t=0}^{T}$ is a martingale, and in satisfies $\mathbb{E}[M_{T}]=1$. By Markov’s inequality, for any $z>0$,
$\Pr(S\geq z)=\Pr(e^{\lambda S}\geq e^{\lambda z})\leq e^{-\lambda z}\mathbb{E}[e^{\lambda S}]$ $=e^{-\lambda z}\underset{=\mathbb{E}[M_{T}]=1}{\underbrace{\mathbb{E}\left[e^{\lambda S}\prod_{t=1}^{T}\sqrt{1-2\lambda\sigma_{t}^{2}}\right]}}\cdot\left(\prod_{t=1}^{T}\sqrt{1-2\lambda\sigma_{t}^{2}}\right)^{-1}$
$=\exp\left(-\lambda z-\frac{1}{2}\sum_{t=1}^{T}\log(1-2\lambda\sigma_{t}^{2})\right)~.$ (20)
To simplify the bound, we note that $\lambda\leq 1/4\beta^{2}$ ensures that $2\lambda\sigma_{t}^{2}\leq\frac{1}{2}$, so we can use the numerical inequality $\log(1-a)\leq 2a$ (which holds for $a\in[0,\frac{1}{2}]$) and get that $\frac{1}{2}\sum_{t=1}^{T}\log(1-2\lambda\sigma_{t}^{2})\leq 2\lambda V$. Plugged into Eq. (20), after a change of variable we see that
$\Pr(S\geq V+z)\leq\exp(-\lambda z+\lambda V)~.$
Concretely, setting $\lambda=1/8\beta^{2}$ and $z=\beta^{2}T+8\beta^{2}\log(1/\delta)$ yields
$\Pr\left(\sum_{t=1}^{T}\xi_{t}^{2}\geq 8\beta^{2}(T+\log(1/\delta))\right)$ $\leq\Pr\left(S\geq V+\beta^{2}T+8\beta^{2}\log(1/\delta))\right)$
$\leq\exp\left(-\frac{T}{8}-\log(1/\delta)+\frac{V}{8\beta^{2}}\right)\leq\delta~.$